# Axion Dark Matter Archaeology with Primordial Gravitational Waves

### v1

@@ 3.2 Estimation of signal-to-noise ratio (SNR)
In any experimental setup, noise from several grounds affects the measured data which demands a precise characterization of the noise spectra. Thus, to quantify the prospects of signal detection in a given experiment, the SNR serves as a fundamental diagnostic tool. For any GW experiment, the SNR can be defined as [190, 196]
@@
${\rm SNR}\equiv\sqrt{\tau_{\rm obs}\int_{f_{\rm min}}^{f_{\rm max}}\text{d}f\left(\frac{\Omega_{\rm GW}(f,\{\theta\})h^{2}}{\Omega_{\rm GW}^{\rm noise}(f)h^{2}}\right)^{2}},$ (3.15)
@@
To ensure broad spectral coverage, we consider multiple GW detectors, including SKA, $\mu$-ARES, LISA, BBO, and ET, collectively spanning the frequency range $[10^{-9},10^{3}]$ Hz. The detailed analysis for noise characteristics of these detectors are discussed in Appendix A. For consistency in our analysis, we have adopted a uniform detection threshold of ${\rm SNR}=10$ across all GW experiments to evaluate their sensitivity to potential signals. The relevant detector specifications are summarized in Table-1.
@@
$\mu$-ARES $\left[10^{-7}-1\right]$ Hz $4$ years
LISA $\left[10^{-4}-1\right]$ Hz $4$ years
BBO $\left[10^{-3}-7\right]$ Hz $4$ years
ET $\left[1-10^{3}\right]$ Hz $5$ years
@@
In Figs. 4-8, we have illustrated the SNR in $m_{Q}-\Lambda$ plane, considering $d=6$ and $d=7$, for SKA, $\mu$-ARES, LISA, BBO and ET, respectively. The analysis is conducted for $n_{T}=0.5$ and $r=0.036$ to exhibit optimal sensitivity. The black and white hatched region come from BBN constraints as in section 2.2. In each figure, the black solid lines represents SNR $=10$, with the region below these lines corresponding SNR $>10$. Our findings reveal that for a given detector and specific ($m_{Q},\Lambda$) point, the SNR is generally higher for $d=6$, compared to $d=7$. This trend can be understood from the sensitivity plots which depicts that increasing the dimension strengthen the domination of early matter era, thereby reducing the detection prospect at the detectors (see, for instance, Fig. 3). Additionally, the plots indicate that SNR increases as $\Lambda$ decreases for fixed $m_{Q}$, consistently across both dimensions and all the detectors. However, the dependence on $m_{Q}$ at fixed $\Lambda$ exhibits opposite behaviour. This trend can be understood by remembering that duration of HQD which increases with increasing $\Lambda$ and decreasing $m_{Q}$. Consequently, achieving a reasonable SNR for each detector requires $\Lambda$ to be large when $m_{Q}$ is also large, a result that holds for both decay dimensions. The dark gray-shaded region in the figures corresponds to GW overproduction as determined by Eq. 3.3. Since $d=6$ results in less damping of the GW spectrum compared to $d=7$, the overproduction occurs for larger values of $\Lambda$ for $d=6$. Due to the same reason, SNR is higher for $d=6$ compared to $d=7$, for a particular $m_{Q}$ and $\Lambda$. Additionally, we observed the BBO, ET and $\mu$-ARES can probe larger $\Lambda$-values.
Figure 6: Descriptions of the figure is same as Fig. 4, but for LISA.
@@
Figure 8: Descriptions of the figure is same as Fig. 4, but for ET.
Additionally in Figs. 4-8, we present the contours for the fraction of axion DM ($\Omega_{a}/\Omega_{\rm DM}$) by dark blue dotted ($10\%$) and dashed ($100\%$) lines. The contours correspond to $f_{a}=10^{12}$ GeV and are shown to indicate how a given $f_{a}$ interacts with our figures. This is important when assessing the complementary information from axion haloscopes and interferometry missions. For $d=7$ decay, GW missions are mostly sensitive for $m_{Q}$, while SKA exhibits promising detection potential to probe the axion DM across a large range of $m_{Q}$ for both dimensions. Thus, SKA, $\mu$-ARES, BBO and ET have the potential interesting regions of the post-inflationary axion scenario.
4 Axion Haloscopes and GW detectors
4.1 Current experimental landscape
As shown in Fig. 2 and the surrounding discussion, the axion haloscope experiments CAPP and ADMX currently constrain small portions of the axion dark matter parameter space. Additionally, in section 3, we presented two existing experimental constraints that limit the parameter space for our axion models when assuming the optimistic inflationary GW background, $r=0.036$ and $n_{T}=0.5$. These are the results from Planck 2018 which limits the GW contribution to $N_{\rm eff}$ and LIGO which constrains $\Omega_{\rm GW}$ in the Hz frequency range. In Fig. 9 we present the additional GW bounds on the parameter space shown in Fig. 2. As exhibited in Figs. 4-8, GW probes are insensitive to the breaking scale $f_{a}$ as well as whether or not the misalignment mechanism produce all to dark matter. However, by fixing $\Omega_{a}=\Omega_{\rm DM}$ we can determine, for a given point on the $\left(m_{Q},\Lambda\right)$-plane, what value for $f_{a}$ is required. This in turn enables us to interpret whether a given axion haloscope experiment is sensitive to this point.
Figure 9: The same as Fig. 2 but with the limits from GWs assuming $n_{T}=0.5$ and $r=0.036$. We show the limits from LIGO and the indirect bound on $\Omega_{\rm GW}$ coming from $\Delta N_{\rm eff}$. We use the result from Planck 2018 and BAO [92].
4.2 Future experimental prospects
The future GW experiments we focus on in this work are discussed in detail in section 3. The prospects for axion haloscope experiments are similarly impressive. In particular the projections of DMRadio [31], FLASH [29], BabyIAXO-RADES [30], and ADMX [94] are projected to probe substantial portions of the $m_{a}$ range that reproduce the correct relic abundance with post-inflationary PQ breaking. In Fig. 10 we only highlight DMRadio and ADMX (green shaded and red shaded regions respectively) for clarity, the sensitivities of FLASH and BabyIAXO-RADES lie in the boundary area of the two experiments shown. Overlaid we show dashed lines which present the sensitivity projections for ET (blue), LISA (magenta), $\mu$-ARES (red), SKA (green), BBO (indigo) as presented in section 3. We also show the projected sensitivities coming from improved measurements of the CMB, which will give a constraint on the additional relativistic degrees of freedom, $\Delta N_{\rm eff}$. The most sensitive of which is reported to be CMB-HD [138] (gray dashed line). We remind the reader that the SNR of each GW experiment below the SNR$=10$ line, therefore each experiment will be sensitive to parameters below the dashed lines.
Figure 10: Future experimental landscape for axion dark matter and GW detectors. In the colored shaded regions we show the parameter regions where ADMX and DMradio a projected to be sensitive to. The dashed lines show the projected reach of future GW detectors assuming $n_{T}=0.5$ and $r=0.036$.
From Fig. 10 one can see that the DMRadio sensitivity region is closer to the $T_{\rm decay}<T_{\rm BBN}$ region, this is because DMRadio will be sensitive to lower values of $m_{a}$ (high $f_{a}$). Therefore, the misalignment production needs to be suppressed to a greater extent, which can be achieved by minimizing $T_{\rm decay}$. A later $T_{\rm decay}$ corresponds to a suppression of $\Omega_{\rm GW}$ at lower $k$ as seen in Fig. 3. The wave-number associated with $T_{\rm BBN}$ is actually lower than the SKA frequency band, therefore for such low $m_{a}$ regions that DMRadio will be sensitive to, SKA is not able to probe. Alternatively, BBO, ET and LISA will be able to probe the DMRadio regions to a greater extent. This is because shorter HQD periods allow for the blue-tilted GW spectra to reach high enough amplitudes in higher frequency bands.
We emphasise that we have chosen the parameters $n_{T}$ and $r$ in order to see the potential GW experiments have to probe high-scale axion models. Therefore the GW lines are not to be read as future constraints on the axion model, they are instead maximum sensitivity lines. In that spirit, we take all the values in the $\left(m_{Q},\Lambda\right)$-plane with ${\rm SNR}\geq 10$ for each experiment. From these points we determine the required $f_{a}$ to reproduce $\Omega_{a}=\Omega_{\rm DM}$. We take the maximum $f_{a}$ obtained and plot the result as a vertical line shown in Fig. 11. We show the $\left(m_{a},g_{a\gamma}\right)$-plane in order to make a stronger connection with axion experiments. We see that BBO, ET have the greatest potential, where the have maximum sensitivity to $f_{a}\sim 2\times 10^{14}\,{\rm GeV}$ and $\mu$-ARES is not far behind $f_{a}\sim 7\times 10^{13}\,{\rm GeV}$, whereas SKA has the potential to probe below $f_{a}\approx 2\times 10^{13}\,{\rm GeV}$ and $f_{a}\approx 2\times 10^{13}\,{\rm GeV}$ for $d=6$. For models where quark decay proceeds via a $d=7$ effective operator, the maximum $f_{a}$ degrades slightly for ET, BBO and $\mu-$ARES but improves slightly for SKA. In both case LISA is essentially only sensitive to scenarios where HQD doesn’t occur or when it doesn’t alter the misalignment mechanism, $T_{\rm decay}>T_{\rm osc}$.
5 Discussion and Conclusion
We propose using the stochastic GW background generated from inflation as a way to determine whether a period of HQD occurred. By using such GWs to trace the early epochs of the Universe, we can obtain information on the particle physics that leads to matter domination. To assess the maximum sensitivity these probes will have, we optimistically take a blue-titled GW power spectrum, $n_{T}=0.5$. Such spectra are well motivated by a range of alternatives to the standard single field, slow-roll inflation. A more comprehensive study would allow the parameters that control the GW power spectrum $A_{T}$ and $n_{T}$, to vary, however by setting optimistic but valid values for such parameters we effectively determine the possible reach of future interferometers.
On the other hand, in the event of a positive detection of a light QCD axion, no $\Delta N_{\rm eff}$ and no of inflationary GWs, HQD would remain as plausible as pre-inflationary case. This is because our investigation has relied on the blue-tilt of the spectrum. It is true that a slightly red-tilted spectrum may still be detectable in future GW detectore. We believe that further investigations into experimental methods to probe the possibility that HQD occurred in the early Universe are warranted. GAN [26], RADES [197], and QUAX [198]. Experiments using a different technology but probing a similar mass range are MADMAX [199], BABYIAXO [30] and FLASH [29]. For lower axion masses in the $10^{-9}$ to $10^{-7}$ eV range, experiments such as SRF [200, 201], DMRadio [31], and ABRACADABRA [202] show particular promise.
Acknowledgement
AC is supported by the National Natural Science Foundation of China (NSFC) through the grants No.12425506, 12375101, 12090060, and 12090064, and the SJTU Double First Class start-up fund (WF220442604). DP thanks Indian Statistical Institute (ISI), Kolkata for financial support through Senior Research Fellowship and the computational facilities of Technology Innovation Hub, ISI Kolkata. The authors also thank Supratik Pal and Lucien Heurtier for fruitful discussions and Leszek Roszkowski for comments of the manuscript.
Appendix A Appendix: Noise modelling for the GW detectors
As already mentioned in Sec. 3.2, estimate of noise spectra associated with the detectors, while assessing the signal. In our study, to asses the signal, we consider the instrumental noise for the aforementioned GW detectors. For LISA and BBO, noise spectra is estimated in terms of effective noise power spectral density ($S(f)$), which can be related with the noise spectra of the detectors as [203]:
$\Omega_{\rm GW}^{\rm noise}(f)\;=\;\left(\frac{4\pi^{2}}{3H_{0}^{2}}\right)f^{3}S(f).$ (A.1)
  • LISA: Instrumental noise of LISA has two key sources: acceleration noise of test mass (acc) and optical metrology noise (omn) [203]. In terms of power spectral density, instrumental noise is expressed as [204]:
    $S_{\rm acc}(f)$ $=9\times 10^{-30}\frac{1}{(2\pi f/1{\rm Hz})^{4}}\left(1+\frac{10^{-4}}{f/1{\rm Hz}}\right)~{}{\rm m^{2}Hz^{-1}},$ (A.2)
    $S_{\rm oms}(f)$ $=\left(1.5\times 10^{-11}{\rm m}\right)^{2}\left[1+\left(\frac{2{\rm mHz}}{f}\right)^{4}\right]\,{\rm Hz}^{-1}.$ (A.3)
    Thus, we have for LISA
    $S_{\rm LISA}(f)$ $=\frac{10}{3L^{2}}\left(2S_{\rm acc}(f)\left(1+{\rm cos}^{2}\left(\frac{f}{c/(2\pi L)}\right)\right)+S_{\rm oms}(f)\right)\left[1+\frac{6}{10}\left(\frac{f}{c/(2\pi L)}\right)^{2}\right],$ (A.4)
    where $L=2.5\times 10^{9}$ m is the arm-length of LISA. Thus, using Eq. (A.1), the GW energy density power spectrum can be calculated.
  • BBO: For BBO, the non sky-averaged instrumental noise power spectral density is [168]
    $S_{\rm BBO}(f)$ $=\left[1.8\times 10^{-49}(f/1{\rm Hz})^{2}+2.9\times 10^{-49}+9.2\times 10^{-52}(f/1{\rm Hz})^{-4}\right]~{}{\rm Hz^{-1}}.$ (A.5)
    Hence, using Eq. (A.1), $\Omega_{\rm GW}^{\rm noise}$ for BBO can be calculated.
For ET and SKA, we have followed Ref. [205] to generate the noise spectra. The characteristic strain of noise ($h_{c}^{\rm noise}$) for the instrumentation of $\mathbf{\mu}$-ARES is provided in Ref. [176]. The noise spectra can be calculated using $h_{c}$ as [132]
$h_{c}^{\rm noise}(f)\simeq 1.26\times^{-18}\left(\frac{1\,{\rm Hz}}{f}\right)\sqrt{h^{2}\Omega^{\rm noise}_{\rm GW}(f)}.$ (A.6)

### v2

@@ III.2 Estimation of signal-to-noise ratio (SNR)
In any experimental setup, noise from several grounds affects the measured data which demands a precise characterization of the noise spectra. Thus, to quantify the prospects of signal detection in a given experiment, the SNR serves as a fundamental diagnostic tool. For any GW experiment, the SNR can be defined as Thrane and Romano (2013); Caprini et al. (2016)
@@
${\rm SNR}\equiv\sqrt{\tau_{\rm obs}\int_{f_{\rm min}}^{f_{\rm max}}\text{d}f\left(\frac{\Omega_{\rm GW}(f,\{\theta\})h^{2}}{\Omega_{\rm GW}^{\rm noise}(f)h^{2}}\right)^{2}},$ (28)
@@
To ensure broad spectral coverage, we consider multiple GW detectors, including SKA, $\mu$-ARES, LISA, BBO, and ET, collectively spanning the frequency range $10^{-9},10^{3}]$ Hz. The detailed analysis for noise characteristics of these detectors are discussed in Appendix A. For consistency in our analysis, we have adopted a uniform detection threshold of ${\rm SNR}=10$ across all GW experiments to evaluate their sensitivity to potential signals. The relevant detector specifications are summarized in Table-1.
In Fig. 4, we show the regions of the $(n_{T},r)$ parameter space where the SNR >10>10 for a range of proposed GW experiments both for absence and presence of HQD. Throughout this analysis, we assume a reheating temperature of $T_{\rm RH}=10^{16}$ GeV and use the same parameter as shown by the black solid line in Fig. 3, i.e. $m_{Q}=10^{14}$ GeV, $d=6$ and $\Lambda=m_{\rm Pl}$, to present the HQD. We also mark $(n_{T},r)=$ $(0,0.036)$ and $(0.5,0.036)$ by purple circle and star in the right panel of the figure which are considered in Fig. 3. Each solid coloured line represents SNR =10=10 contour, corresponding to each GW detector, with the right hand portion of each contour indicates SNR >10>10. Although the SNR enhances as $n_{T}$ increases, higher $n_{T}$ values are excluded by Planck-18+BAO data, indicated by grey hatched region. However, compared to standard scenario (without HQD), presence of HQD allows higher values of $n_{T}$ (discussed in the previous section) as also shown the plots. The vertical dashed line at $n_{T}\simeq-r/8$ denotes the slow-roll consistency relation, which can be tested by BBO, ET, and $\mu$-ARES. In contrast, the detection prospects for LISA and SKA require a blue-tilted spectrum with $n_{T}>0$. Overall, the figure highlights that primordial GWs can be probed across a broad range of parameters. While enhanced detectability favours positive values of $n_{T}$, the slow-roll prediction remains encompassed within this parameter space. Moreover, any such signal is expected to be further modified by an early matter domination phase, driven by the presence of heavy quark, which constitutes the primary focus of our investigation.
Figure 4: SNR in the $n_{T}-r$ plane across several GW detectors. Left panel corresponds to the standard scenario (without HQD) whereas right panel presents the HQD scenario for the same set of benchmark parameters ($m_{Q}=10^{14}$ GeV, $d=6$ and $\Lambda=m_{\rm Pl}$) as in Fig. 3. The purple star and circle in the right panel mark the benchmark point in the $n_{T}-r$ plane, which are also utilized in Fig. 3. Both panels assume $T_{\rm RH}=10^{16}$ GeV. Coloured lines correspond to $\rm SNR=10$ contour for the respective detector, with the region to the right $\rm SNR>10$. Hatched region presents the excluded region by Planck-18+BAO due to overproduction GWs.
@@
$\mu$-ARES $\left[10^{-7}-1\right]$ Hz 44 years
LISA $\left[10^{-4}-1\right]$ Hz 44 years
BBO $\left[10^{-3}-7\right]$ Hz 44 years
ET $\left[1-10^{3}\right]$ Hz 55 years
@@
In Figs. 5-6, we have illustrated the SNR in $m_{Q}-\Lambda$ plane, considering $d=6$ and $d=7$, for LISA and ET, respectively (SNR plots for the other GW detectors are illustrated in the appendix B). The analysis is conducted for $n_{T}=0.5$ and $r=0.036$ to exhibit optimal sensitivity. The black and white hatched region come from BBN constraints as in section II.2. In each figure, the black solid lines represents SNR =10=10, with the region below these lines corresponding SNR >10>10. Our findings reveal that for a given detector and specific ($m_{Q},\Lambda$) point, the SNR is generally higher for $d=6$, compared to $d=7$. This trend can be understood from the sensitivity plots which depicts that increasing the dimension strengthen the domination of early matter era, thereby reducing the detection prospect at the detectors (see, for instance, Fig. 3). Additionally, the plots indicate that SNR increases as $\Lambda$ decreases for fixed $m_{Q}$, consistently across both dimensions and all the detectors. However, the dependence on $m_{Q}$ at fixed $\Lambda$ exhibits opposite behaviour. This trend can be understood by remembering that duration of HQD which increases with increasing $\Lambda$ and decreasing $m_{Q}$. Consequently, achieving a reasonable SNR for each detector requires $\Lambda$ to be large when $m_{Q}$ is also large, a result that holds for both decay dimensions. The dark gray-shaded region in the figures corresponds to GW overproduction as determined by Eq. (16). Since $d=6$ results in less damping of the GW spectrum compared to $d=7$, the overproduction occurs for larger values of $\Lambda$ for $d=6$. Due to the same reason, SNR is higher for $d=6$ compared to $d=7$, for a particular $m_{Q}$ and $\Lambda$. Additionally, we observed the BBO, ET and $\mu$-ARES can probe larger $\Lambda$-values.
Figure 5: Illustration of SNR in the $m_{Q}-\Lambda$ plane for LISA across different dimensions. Black solid line represents ${\rm SNR}=10$, with the region below this line corresponding to ${\rm SNR}>10$, as indicated by the colour scale. The gray-shaded region denotes the parameter space excluded by BBN constraints due to the overproduction of gravitational waves. Both the dimensions are for $n_{T}=0.5$ and $r=0.036$.
Figure 6: Descriptions of the figure is same as Fig. 5, but for ET.
Additionally in the figures, we present the contours for the fraction of axion DM ($\Omega_{a}/\Omega_{\rm DM}$) by dark blue dotted ($10\%$) and dashed ($100\%$) lines. The contours correspond to $f_{a}=10^{12}$ GeV and are shown to indicate how a given $f_{a}$ interacts with our figures. This is important when assessing the complementary information from axion haloscopes and interferometry missions. For $d=7$ decay, GW missions are mostly sensitive for $m_{Q}$, while ET exhibits promising detection potential to probe the axion DM across a large range of DM mass for both dimensions.
IV Axion Haloscopes and GW detectors
IV.1 Current experimental landscape
As shown in Fig. 2 and the surrounding discussion, the axion haloscope experiments CAPP and ADMX currently constrain small portions of the axion dark matter parameter space. Additionally, in section III, we presented two existing experimental constraints that limit the parameter space for our axion models when assuming the optimistic inflationary GW background, $r=0.036$ and $n_{T}=0.5$. These are the results from Planck 2018 which limits the GW contribution to $N_{\rm eff}$ and LIGO which constrains $\Omega_{\rm GW}$ in the Hz frequency range. In Fig. 7 we present the additional GW bounds on the parameter space shown in Fig. 2. As exhibited in Figs. 10-6, GW probes are insensitive to the breaking scale $f_{a}$ as well as whether or not the misalignment mechanism produce all to dark matter. However, by fixing $\Omega_{a}=\Omega_{\rm DM}$ we can determine, for a given point on the $\left(m_{Q},\Lambda\right)$-plane, what value for $f_{a}$ is required. This in turn enables us to interpret whether a given axion haloscope experiment is sensitive to this point.
Figure 7: The same as Fig. 2 but with the limits from GWs assuming $n_{T}=0.5$ and $r=0.036$. We show the limits from LIGO and the indirect bound on $\Omega_{\rm GW}$ coming from $\Delta N_{\rm eff}$. We use the result from Planck 2018 and BAO Aghanim et al. (2020).
IV.2 Future experimental prospects
The future GW experiments we focus on in this work are discussed in detail in section III. The prospects for axion haloscope experiments are similarly impressive. In particular the projections of DMRadio Brouwer et al. (2022), FLASH Alesini et al. (2023), BabyIAXO-RADES Ahyoune et al. (2023), and ADMX Stern (2016) are projected to probe substantial portions of the $m_{a}$ range that reproduce the correct relic abundance with post-inflationary PQ breaking. In Fig. 8 we only highlight DMRadio and ADMX (green shaded and red shaded regions respectively) for clarity, the sensitivities of FLASH and BabyIAXO-RADES lie in the boundary area of the two experiments shown. Overlaid we show dashed lines which present the sensitivity projections for ET (blue), LISA (magenta), $\mu$-ARES (red), SKA (green), BBO (indigo) as presented in section III. We also show the projected sensitivities coming from improved measurements of the CMB, which will give a constraint on the additional relativistic degrees of freedom, $\Delta N_{\rm eff}$. The most sensitive of which is reported to be CMB-HD Aiola et al. (2022) (gray dashed line). We remind the reader that the SNR of each GW experiment below the SNR=10=10 line, therefore each experiment will be sensitive to parameters below the dashed lines.
Figure 8: Future experimental landscape for axion dark matter and GW detectors. In the colored shaded regions we show the parameter regions where ADMX and DMradio are projected to be sensitive to. The dashed lines show the projected reach of future GW detectors assuming $n_{T}=0.5$ and $r=0.036$.
From Fig. 8 one can see that the DMRadio sensitivity region is closer to the $T_{\rm dec}<T_{\rm BBN}$ region, this is because DMRadio will be sensitive to lower values of $m_{a}$ (high $f_{a}$). Therefore, the misalignment production needs to be suppressed to a greater extent, which can be achieved by minimizing $T_{\rm dec}$. A later $T_{\rm dec}$ corresponds to a suppression of $\Omega_{\rm GW}$ at lower $k$ as seen in Fig. 3. The wave-number associated with $T_{\rm BBN}$ is actually lower than the SKA frequency band, therefore for such low $m_{a}$ regions that DMRadio will be sensitive to, SKA is not able to probe. Alternatively, BBO, ET and LISA will be able to probe the DMRadio regions to a greater extent. This is because shorter HQD periods allow for the blue-tilted GW spectra to reach high enough amplitudes in higher frequency bands.
We emphasise that we have chosen the parameters $n_{T}$ and $r$ in order to see the potential GW experiments have to probe high-scale axion models. Therefore the GW lines are not to be read as future constraints on the axion model, they are instead maximum sensitivity lines. In that spirit, we take all the values in the $\left(m_{Q},\Lambda\right)$-plane with ${\rm SNR}\geq 10$ for each experiment. From these points we determine the required $f_{a}$ to reproduce $\Omega_{a}=\Omega_{\rm DM}$. We take the maximum $f_{a}$ obtained and plot the result as a vertical line shown in Fig. 9. We show the $\left(m_{a},g_{a\gamma}\right)$-plane in order to make a stronger connection with axion experiments. We see that BBO, ET have the greatest potential, where the have maximum sensitivity to $f_{a}\sim 2\times 10^{14}\,{\rm GeV}$ and $\mu$-ARES is not far behind $f_{a}\sim 7\times 10^{13}\,{\rm GeV}$, whereas SKA has the potential to probe below $f_{a}\approx 2\times 10^{13}\,{\rm GeV}$ and $f_{a}\approx 2\times 10^{13}\,{\rm GeV}$ for $d=6$. For models where quark decay proceeds via a $d=7$ effective operator, the maximum $f_{a}$ degrades slightly for ET, BBO and $\mu$-ARES but improves slightly for SKA. In both case LISA is essentially only sensitive to scenarios where HQD doesn’t occur or when it doesn’t alter the misalignment mechanism, $T_{\rm dec}>T_{\rm osc}$.
V Discussion and Conclusion
We propose using the stochastic GW background generated from inflation as a way to determine whether a period of HQD occurred. By using such GWs to trace the early epochs of the Universe, we can obtain information on the particle physics that leads to matter domination. We investigated $n_{T}=0$ spectrum and its detectability with the GW detectors and then to assess the maximum sensitivity these probes will have, we optimistically take a blue-titled GW power spectrum, $n_{T}=0.5$ in Fig. 3 for illustration purpose only. Such spectra are well motivated by a range of alternatives to the standard single field, slow-roll inflation, to list a few of them Liddle and Lyth (1993b); Brandenberger et al. (2007); Calcagni et al. (2014); Fujita et al. (2019); Cook and Sorbo (2012); Mukohyama et al. (2014); Anber and Sorbo (2010); Baldi et al. (2005); Oikonomou (2021); Biagetti et al. (2013); Caldwell and Devulder (2018); Dimastrogiovanni et al. (2018); Adshead and Sfakianakis (2017); Kuroyanagi et al. (2021); Afzal et al. (2023). A more comprehensive study would allow the parameters that control the GW power spectrum $A_{T}$ and $n_{T}$, to vary, however by setting optimistic but valid values for such parameters we effectively determine the possible reach of future interferometers.
On the other hand, in the event of a positive detection of a light QCD axion, no $\Delta N_{\rm eff}$ and no of inflationary GWs, HQD would remain as plausible as pre-inflationary case. This is because our investigation has relied on the blue-tilt of the spectrum. It is true that a slightly red-tilted spectrum may still be detectable in future GW detectore. We believe that further investigations into experimental methods to probe the possibility that HQD occurred in the early Universe are warranted. GAN McAllister et al. (2017), RADES Melcón et al. (2018), and QUAX Alesini et al. (2021). Experiments using a different technology but probing a similar mass range are MADMAX Brun et al. (2019), BABYIAXO Ahyoune et al. (2023) and FLASH Alesini et al. (2023). For lower axion masses in the $10^{-9}$ to $10^{-7}$ eV range, experiments such as SRF Berlin et al. (2021); Tang et al. (2024), DMRadio Brouwer et al. (2022), and ABRACADABRA Salemi et al. (2021) show particular promise.
Acknowledgement
AC is supported by the National Natural Science Foundation of China (NSFC) through the grants No.12425506, 12375101, 12090060, and 12090064, and the SJTU Double First Class start-up fund (WF220442604). DP thanks Indian Statistical Institute (ISI), Kolkata for financial support through Senior Research Fellowship and the computational facilities of Technology Innovation Hub, ISI Kolkata. The authors also thank Supratik Pal, Lucien Heurtier and Enrico Nardi for fruitful discussions and Leszek Roszkowski for comments of the manuscript.
Appendix A Noise modelling for the GW detectors
As already mentioned in Sec. III.2, estimate of noise spectra associated with the detectors, while assessing the signal. In our study, to asses the signal, we consider the instrumental noise for the aforementioned GW detectors. For LISA and BBO, noise spectra is estimated in terms of effective noise power spectral density ($S(f)$), which can be related with the noise spectra of the detectors as Gowling and Hindmarsh (2021):
$\Omega_{\rm GW}^{\rm noise}(f)\;=\;\left(\frac{4\pi^{2}}{3H_{0}^{2}}\right)f^{3}S(f).$ (29)
  • LISA: Instrumental noise of LISA has two key sources: acceleration noise of test mass (acc) and optical metrology noise (omn) Gowling and Hindmarsh (2021). In terms of power spectral density, instrumental noise is expressed as Breitbach et al. (2019):
    $S_{\rm acc}(f)$ $=9\times 10^{-30}\frac{1}{(2\pi f/1{\rm Hz})^{4}}\left(1+\frac{10^{-4}}{f/1{\rm Hz}}\right)~{\rm m^{2}Hz^{-1}},$ (30)
    $S_{\rm oms}(f)$ $=\left(1.5\times 10^{-11}{\rm m}\right)^{2}\left[1+\left(\frac{2{\rm mHz}}{f}\right)^{4}\right]\,{\rm Hz}^{-1}.$ (31)
    Thus, we have for LISA
    $S_{\rm LISA}(f)$ $=\frac{10}{3L^{2}}\left(2S_{\rm acc}(f)\left(1+{\rm cos}^{2}\left(\frac{f}{c/(2\pi L)}\right)\right)\right$. (32)
    $\left.+S_{\rm oms}(f)\right)\left[1+\frac{6}{10}\left(\frac{f}{c/(2\pi L)}\right)^{2}\right],$
    where $L=2.5\times 10^{9}$ m is the arm-length of LISA. Thus, using Eq. (29), the GW energy density power spectrum can be calculated.
  • BBO: For BBO, the non sky-averaged instrumental noise power spectral density is Yagi et al. (2011)
    $S_{\rm BBO}(f)$ $=\left[1.8\times 10^{-49}(f/1{\rm Hz})^{2}+2.9\times 10^{-49}\right$. (33)
    $\left.+9.2\times 10^{-52}(f/1{\rm Hz})^{-4}\right]~{\rm Hz^{-1}}.$
    Hence, using Eq. (29), $\Omega_{\rm GW}^{\rm noise}$ for BBO can be calculated.
For ET and SKA, we have followed Ref. Schmitz (2021) to generate the noise spectra. The characteristic strain of noise ($h_{c}^{\rm noise}$) for the instrumentation of $\mathbf{\mu}$-ARES is provided in Ref. Sesana et al. (2021). The noise spectra can be calculated using $h_{c}$ as Maggiore (2000)
$h_{c}^{\rm noise}(f)\simeq 1.26\times^{-18}\left(\frac{1\,{\rm Hz}}{f}\right)\sqrt{h^{2}\Omega^{\rm noise}_{\rm GW}(f)}.$ (34)
Appendix B Accessing the GW signal for SKA, $\mu$-ARES and BBO
Figs. 10-12 show the SNR in $m_{Q}-\Lambda$ plane for $d=6$ and $d=7$, respectively for SKA, $\mu$-ARES and BBO. As in the LISA and ET analysis in Sec. III.2, the results are presented for $n_{T}=0.5$ and $r=0.36$, considering instrumental noise of each detector. In all figures, the black solid curve denotes SNR=10=10, with the region below this curve corresponding to SNR>10>10. As discussed in Sec. III.2, the case $d=6$ yields larger SNR than $d=7$ across all the GW detectors. Moreover, the SNR is directly proportional to $m_{Q}$ for a fixed $\Lambda$, however, it is inversely proportional to $\Lambda$ for a fixed $m_{Q}$ (discussed in detailed in Sec III.2). The dark blue dotted and dashed curves represent the axion DM fraction $10\%$ and $100\%$, respectively. Thus, the detectors exhibit promising regions in the $m_{Q}-\Lambda$ plane where axion DM can probed over a broad parameter range.